1/5+1/7=?
1/6+1/8=?
16/9 is equal to?

Answer by rfer(16322)

1/5+1/7=
7/35+5/35=12/35
---------------
1/6+1/8=
4/24+3/24=7/24
----------------
16/9=1 7/9